# Multi-Agent Collaborative Filtering: Orchestrating Users and Items for Agentic Recommendations

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems are shifting from static predictors toward interactive agents. Recent agentic recommenders in web applications, such as online shopping assistants and movie finders, cast the system as a large language model (LLM) agent that can plan, reason, call tools, and interact with users (Yao et al., 2023). These systems usually follow one of two patterns. The first uses a single LLM agent in a plan–and–execute loop (Wang et al., 2024a; Huang et al., 2025) that reasons step-wise and calls retrieval or recommendation tools but still acts as a single decision-maker, making it hard to draw on diverse preference signals. The second organizes multiple LLM agents into task-decomposition pipelines with generic dialogue or planning roles (Wang et al., 2024b; Fang et al., 2024; Yu et al., 2025). These multi-agent designs introduce role specialization and coordination , but their roles are loosely tied to preference signal sources such as similar users or relevant items. Thus, while these approaches improve interactivity and tool use, they lack an effective way to bring collaborative signals into the agentic recommendation process.
As the foundation for many modern recommender systems, collaborative filtering (CF) has long shown that similar users and relevant items hold strong predictive value (Resnick et al., 1994; Sarwar et al., 2001). Neural CF models (He et al., 2017) further capture such signals with learned representations. Recent AgentCF (Zhang et al., 2024) extends this idea by modeling users and items as LLM agents that interact under logged user–item histories for behavior simulation. Despite these advances, most LLM-based agentic recommenders still treat user–item histories only as context in prompts and do not structure or combine neighborhood evidence as CF does. As a result, they struggle to take advantage of the collaborative signals during recommendation, especially when similar users and correlated items provide complementary information.
@@ 2.1. Problem Definition
  • $\mathrm{RetrieveByItem}(i,k)$ retrieves $k$ items whose embeddings are most similar to a given item ii.
@@ 3.1. Experimental Setup
We evaluate our method on the three Amazon datasets: Clothing, Beauty, and Music (McAuley et al., 2015). We adopt the same user simulator-driven evaluation protocol as in (Yu et al., 2025) to measure the recommendation quality. H@10 (HitRatio@10) measures how often relevant items appear in the top-1010 positions of the recommendation list. N@10 (NDCG@10) evaluates the ranking quality of top-1010 items in the recommendation list. We adopt the same data splits and natural language query sets as in Yu et al. (2025).

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792931isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems are shifting from static predictors toward interactive agents. Recent agentic recommenders in web applications, such as online shopping assistants and movie finders, cast the system as a large language model (LLM) agent that can plan, reason, call tools, and interact with users (Huang et al., 2025a; Xia et al., 2025a). These systems usually follow one of two patterns. The first uses a single LLM agent in a plan–and–execute loop (Xia et al., 2025b; Huang et al., 2025b) that reasons step-wise and calls retrieval or recommendation tools but still acts as a single decision-maker, making it hard to draw on diverse preference signals (Shen et al., 2025). The second organizes multiple LLM agents into task-decomposition pipelines with generic dialogue or planning roles (Wang et al., 2024; Fang et al., 2024; Yu et al., 2025). These multi-agent designs introduce role specialization and coordination, but their roles are loosely tied to preference signal sources such as similar users or relevant items. Thus, while these approaches improve interactivity and tool use, they lack an effective way to bring collaborative signals into the agentic recommendation process.
As the foundation for many modern recommender systems, collaborative filtering (CF) has long shown that similar users and relevant items hold strong predictive value (Resnick et al., 1994; Sarwar et al., 2001; Xia et al., 2023). Recent AgentCF (Zhang et al., 2024) extends this idea by modeling users and items as LLM agents that interact under logged user–item histories for behavior simulation. Despite these advances, most LLM-based agentic recommenders still treat user–item histories only as context in prompts and do not structure or combine neighborhood evidence as CF does. As a result, they struggle to take advantage of the collaborative signals during recommendation, especially when similar users and correlated items provide complementary information.
@@ 2.1. Problem Definition
  • $\mathrm{RetrieveByItem}(i,k)$ retrieves $k$ items whose embeddings are most similar to a given item $i$.
@@ 3.1. Experimental Setup
We evaluate our method on the three Amazon datasets: Clothing, Beauty, and Music (McAuley et al., 2015). We adopt the same user simulator-driven evaluation protocol as in (Yu et al., 2025) to measure the recommendation quality. H@10 (HitRatio@10) measures how often relevant items appear in the top-$10$ positions of the recommendation list. N@10 (NDCG@10) evaluates the ranking quality of top-$10$ items in the recommendation list. We adopt the same data splits and natural language query sets as in Yu et al. (2025).